# Model-independent predictions for decays of double-heavy hadrons into pairs of heavy hadrons

### v2

@@ Abstract
Double-heavy hadrons can decay into pairs of heavy hadrons through transitions from confining Born-Oppenheimer potentials to hadron-pair potentials with the same Born-Oppenheimer quantum numbers. The transitions are also constrained by conservation of angular momentum and parity. From these constraints, we derive model-independent selection rules for decays of double-heavy hadrons into pairs of heavy hadrons. The coupling potentials are expressed as sums of products of Born-Oppenheimer transition amplitudes and angular-momentum coefficients. If there is a single dominant Born-Oppenheimer transition amplitude, it factors out of the coupling potentials between double-heavy hadrons in the same Born-Oppenheimer multiplet and pairs of heavy hadrons in specific heavy-quark-spin-symmetry doublets. If furthermore the kinetic energies of the heavy hadrons are much larger than their spin splittings, we obtain analytic expressions for the relative partial decay rates in terms of Wigner 6-$j$ and 9-$j$ symbols. We consider in detail the decays of quarkonia and quarkonium hybrids into the lightest heavy-meson pairs. For quarkonia, our model-independent selection rules and relative partial decay rates agree with previous results from quark-pair-creation models in simple cases and give stronger results in other cases. For quarkonium hybrids, we find disagreement even in simple cases.
@@ I Introduction
The field of double-heavy hadron spectroscopy began in 1974 with the discovery of the $J\!/\!\psi$ meson [1, 2]. By the end of the century, dozens of double-heavy mesons had been discovered, all of which matched the pattern of quarkonium energy levels predicted by quark models. The spectrum of double-heavy hadrons was generally believed to be well understood. But all this changed in 2003 with the discovery of the $X(3872)$ meson [3]. Its observed decays revealed that it is a tetraquark meson whose constituents consist of a charm-anticharm pair and an additional light quark-antiquark pair. Since then, dozens of exotic double-heavy hadrons have been discovered, inaugurating a renaissance of double-heavy-hadron spectroscopy; see, for instance, the reviews in Refs. [4, 5].
@@
In this paper, we focus on decays of double-heavy hadrons into pairs of heavy hadrons. These strong decays are subject to the low energy, nonperturbative regime of QCD, which is notoriously difficult to access from first principles. Hadron resonances with 2-body (and some 3-body) strong decay channels can be calculated ab initio using lattice QCD and the Lüscher formalism; see Ref. [6] for a review. But double-heavy hadrons that are above the threshold for a heavy-hadron pair have many multibody strong decay channels, including lower-energy double-heavy hadrons plus light hadrons. So, first-principles calculations of double-heavy hadron resonances using the Lüscher formalism might be impractical.
@@
An alternative approach to the study of double-heavy hadrons from first principles is the Born-Oppenheimer (B-O) approximation for QCD [7]. The B-O approximation is based on the assumption that gluons and light quarks respond almost instantaneously to the motion of the heavy (anti)quarks, since the latter have a mass $m_{Q}$ that is much larger than the nonperturbative energy scale of QCD $\Lambda_{\textup{QCD}}$. In this approximation, the energy levels of the gluon and light-quark fields in the presence of two static color sources, which can be calculated using lattice QCD, play the role of B-O potentials that determine the motion of the heavy (anti)quarks through a Schrödinger equation. It has been shown that the B-O approximation can be formulated as a rigorous effective theory of QCD [8, 9, 10, 11].
@@
There are two qualitatively different kinds of B-O potentials in QCD. The first kind consists of confining potentials that increase linearly at large distances [12]. The solutions to the Schrödinger equation in these potentials are bound states associated with compact double-heavy hadrons. The second kind consists of potentials that at large distances approach the scattering threshold of a pair of heavy hadrons [13]. The solutions to the Schrödinger equation in these potentials include scattering states of heavy-hadron pairs. They might also include bound states associated with heavy-hadron molecules.
@@
The B-O potentials for double-heavy hadrons and for heavy-hadron pairs mix with each other [13, 14]. Hence, any double-heavy hadron is associated with a pole of the S-matrix for heavy-hadron pairs. For double-heavy hadrons that are stable against decays into heavy-hadron pairs, the imaginary part of the pole is proportional to the total decay width into lower-energy double-heavy hadrons plus light hadrons. For double-heavy hadrons associated with heavy-hadron-pair scattering resonances, the imaginary part of the pole has additional contributions from decays into heavy-hadron pairs.
@@
One way to calculate partial decay widths of double-heavy-hadrons into heavy-hadron pairs using the B-O approximation for QCD is to calculate self-energy corrections to “bare” double-heavy-hadron states by resummation of heavy-hadron loop diagrams, as was done in Refs. [15, 16]. But it should be noted that this approach can lead to inconsistent results for double-heavy hadrons that are strongly coupled to a nearby heavy-hadron-pair threshold [15]. Another way is to calculate resonances directly from the Schrödinger equation with coupled double-heavy-hadron and heavy-hadron-pair channels, as was done in Refs. [17, 18, 19]. This approach should be preferred as it respects unitarity and it yields consistent results for resonances that are close to a heavy-hadron-pair threshold [18].
@@
The remainder of this paper is organized as follows. In Section II, we define static hadrons and heavy hadrons. We then construct heavy-hadron-pair states and discuss their symmetries. In Section III, we derive general expressions for the coupling potentials between double-heavy hadrons and heavy-hadron pairs as sums of products of B-O transition amplitudes and angular-momentum coefficients. From these expressions, we derive model-independent selection rules for decays of double-heavy hadrons. In some cases, we also obtain analytic predictions for relative partial decay rates. In Section IV, we discuss in detail the decays of quarkonia and quarkonium hybrids into pairs of heavy mesons. In Section V, we compare our results with previous ones from quark-pair-creation models. Finally, we summarize these results in Section VI.
@@ II.1 Single Heavy Hadron
$\mathcal{P}\biggl{|}\begin{matrix}\frac{1}{2}^{\pm}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ =±$=\pm\biggl{|}\begin{matrix}\frac{1}{2}^{\pm}\\[-5.0pt] m\end{matrix}\biggr{\rangle},$ (1e)
$\mathcal{C}\biggl{|}\begin{matrix}\frac{1}{2}^{\pm}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ $=\biggl{|}\begin{matrix}\frac{1}{2}^{\mp}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (1j)
@@
$\mathcal{P}\biggl{|}\begin{matrix}j^{\pi}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ $=\pi\biggl{|}\begin{matrix}j^{\pi}\\[-5.0pt] m\end{matrix}\biggr{\rangle},$ (2e)
$\mathcal{C}\biggl{|}\begin{matrix}j^{\pi}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ $=\biggl{|}\begin{matrix}j^{\pi(-1)^{2j}}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (2j)
Note that a static hadron and its charge conjugate have the same (opposite) parity if $j$ is an integer (half-integer), hence the multiplicative factor $(-1)^{2j}$ in the parity superscript on the right side of Eq. (2j).
@@
[rb,eb] where $j$ can be $J+\frac{1}{2}$ or $\bigl{\lvert}J-\frac{1}{2}\bigr{\rvert}$ and $\bigl{\langle}\begin{smallmatrix}j_{1}&j_{2}\\ m_{1}&m_{2}\\ \end{smallmatrix}\big{|}\begin{smallmatrix}j_{3}\\ m_{3}\\ \end{smallmatrix}\bigr{\rangle}$ is a Clebsch-Gordan coefficient. The hadron is a meson if $J$ is an integer and a baryon if $J$ is a half-integer. We have adopted the standard order of writing the spin/parity of the heavy quark or antiquark first and the angular-momentum/parity of the light-QCD fields second. From Eqs. (II.1)-(II.1), the transformations under $\mathcal{P}$ and $\mathcal{C}$ of the heavy-hadron state in Eq. (3) are:
$\mathcal{P}\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j^{\pi}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle}$ =±$=\pm\pi\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j^{\pi}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle},$ (4e)
$\mathcal{C}\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j^{\pi}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle}$ $=\biggl{|}\Bigl{(}\tfrac{1}{2}^{\mp},j^{\pi(-1)^{2j}}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle}.$ (4j)
@@
Since the spin of a heavy quark or antiquark is conserved up to corrections suppressed by $1/m_{Q}$, the mass difference between two heavy hadrons corresponding to the same static hadron $j^{\pi}$ but with different spins $J=\bigl{\lvert}j-\frac{1}{2}\bigr{\rvert}$ and $J=j+\frac{1}{2}$ must also be suppressed by $1/m_{Q}$. Therefore, heavy hadrons form approximately degenerate heavy-quark-spin-symmetry (HQSS) doublets labeled by the light-QCD angular-momentum/parity $j^{\pi}$ and other quantum numbers.
@@ II.2 Pair of Static Hadrons
Let us now consider a pair of static hadrons at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$ with light-QCD quantum numbers $j_{1}^{\pi_{1}}$ and $j_{2}^{\pi_{2}}$, respectively. Since the angular-momentum/parity quantum numbers correspond to rotations and reflections around different points in space, we specify the position ±$\pm\frac{1}{2}\bm{r}$ by an argument (±)plus-or-minus($\pm$)( ± ). The static-hadron-pair states are therefore labeled by $\bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\bigr{)}$.
@@
Light QCD in the presence of two static color sources at $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$ has cylindrical symmetries consisting of rotations around the $\bm{\hat{r}}$ axis and reflections through planes containing $\bm{\hat{r}}$. We denote the reflection corresponding to a specific plane (which need not be specified) by $\mathcal{R}$. There is also an additional discrete symmetry that depends on whether the sources are two static quarks $(\bm{3},\bm{3})$ or one static quark and one static antiquark $(\bm{3},\bm{3}^{\ast})$. [^1] We do not consider the $(\bm{3}^{\ast},\bm{3}^{\ast})$ case, which is just the charge conjugate of $(\bm{3},\bm{3})$. In the presence of $(\bm{3},\bm{3})$ sources, light QCD is symmetric under parity inversion through the origin, $\mathcal{P}$. In the presence of $(\bm{3},\bm{3}^{\ast})$ sources, light QCD is symmetric under the combined transformation $\mathcal{C}\mathcal{P}$, with $\mathcal{C}$ for charge conjugation. For the sake of brevity, we shall refer to this discrete symmetry as $[\mathcal{C}]\mathcal{P}$, which is defined to be $\mathcal{P}$ in the $(\bm{3},\bm{3})$ case and $\mathcal{C}\mathcal{P}$ in the $(\bm{3},\bm{3}^{\ast})$ case. We refer to the group formed by the cylindrical symmetries and the $[\mathcal{C}]\mathcal{P}$ symmetry as the B-O symmetry group.
@@
The B-O symmetries imply that the eigenstates of the light-QCD Hamiltonian can be chosen to be simultaneously eigenstates of $\bm{J}_{\textup{light}}\cdot\bm{\hat{r}}$, that is, the projection of the light-QCD angular momentum onto the axis passing through the sources. We denote its eigenvalues by $\lambda$. They can be also chosen to be simultaneously eigenstates of $[\mathcal{C}]\mathcal{P}$. We denote its eigenvalues by $\eta$. In the sector with $\lambda=0$, the reflection operator $\mathcal{R}$ can also be diagonalized. We denote its eigenvalues by $\epsilon$. We refer to $\lambda$ and $\eta$ and also $\epsilon$ if $\lambda=0$ as B-O quantum numbers. Alternatively, the light-QCD Hamiltonian, $\lvert\bm{J}_{\textup{light}}\cdot\hat{\bm{r}}\rvert$, $[\mathcal{C}]\mathcal{P}$, and $\mathcal{R}$ can be simultaneously diagonalized. Thus an alternative choice for the B-O quantum numbers is $\lvert\lambda\rvert$, $\eta$, and $\epsilon$. It is customary to denote these B-O quantum numbers using the notation $\Lambda_{\eta}^{\epsilon}$, where $\Lambda=\lvert\lambda\rvert$, the subscript $\eta$ is $g$ or $u$ if the $[\mathcal{C}]\mathcal{P}$ eigenvalue $\eta$ is $+1$ or $-1$, and the superscript $\epsilon$ is $+$ or -- if the reflection eigenvalue $\epsilon$ is $+1$ or $-1$. If $\Lambda>0$, the superscript $\epsilon$ is often omitted because cylindrical symmetry requires the $\Lambda_{\eta}^{+}$ and $\Lambda_{\eta}^{-}$ states to be degenerate in energy. It is commonplace to specify integer values of $\Lambda$ with an uppercase Greek letter instead of a number, according to the code $\Lambda\to\Sigma,\Pi,\Delta$ for $\Lambda=0,1,2$ and so on.
@@
The sign is the product of a factor $(-1)^{j_{1}+j_{2}-j^{\prime}}$ from the symmetries of Clebsch-Gordan coefficients and a factor $(-1)^{4j_{1}j_{2}}$ from changing the order of the light-QCD operators. [^2] The factor $(-1)^{4j_{1}j_{2}}$ is $-1$ if both static hadrons are static mesons and $+1$ otherwise. From Eqs. (II.1), the transformations under $\mathcal{P}$ and $\mathcal{C}$ of the static-hadron-pair state in Eq. (5) are:
$\mathcal{P}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}=$ $\pi_{1}\pi_{2}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(-),j_{2}^{\pi_{2}}(+)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle},$ (7e)
$\mathcal{C}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}=$ $\biggl{|}\Bigl{(}j_{1}^{\pi_{1}(-1)^{2j_{1}}}(+),j_{2}^{\pi_{2}(-1)^{2j_{2}}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}.$ (7j)
On the right side of Eq. (7e), the position labels ()(-)( - ) and (+)(+)( + ) can be put into the standard order by using Eq. (6).
There are special cases in which the static-hadron-pair state in Eq. (5) is a simultaneous eigenstate of $[\mathcal{C}]\mathcal{P}$. For $(\bm{3},\bm{3})$ sources, the static-hadron-pair state is a simultaneous eigenstate of $\mathcal{P}$ when the two static hadrons are identical, which requires $j_{2}^{\pi_{2}}=j_{1}^{\pi_{1}}$. After reordering the ()(-)( - ) and (+)(+)( + ) states using Eq. (6), the $\mathcal{P}$ eigenvalue is the product of $\pi_{1}^{2}=1$ and $(-1)^{4j_{1}^{2}+2j_{1}-j^{\prime}}$. For $(\bm{3},\bm{3}^{\ast})$ sources, the static-hadron-pair state is a simultaneous eigenstate of $\mathcal{C}\mathcal{P}$ when the two static hadrons are charge conjugates, which requires $j_{2}^{\pi_{2}}=j_{1}^{\pi_{1}(-1)^{2j_{1}}}$. After reordering the ()(-)( - ) and (+)(+)( + ) states using Eq. (6), the $\mathcal{C}\mathcal{P}$ eigenvalue is the product of $\pi_{1}^{2}(-1)^{2j_{1}}=(-1)^{2j_{1}}$ and $(-1)^{4j_{1}^{2}+2j_{1}-j^{\prime}}$. Since $2j_{1}(2j_{1}+1)$ is even, the $[\mathcal{C}]\mathcal{P}$ eigenvalue reduces to
$\eta=\begin{cases}(-1)^{j^{\prime}}&\text{for $(\bm{3},\bm{3})$ sources,}\\ (-1)^{j^{\prime}+2j_{1}}&\text{for $(\bm{3},\bm{3}^{\ast})$ sources.}\end{cases}$ (8)
Note that $j^{\prime}$ is an integer if the static-hadron-pair state is an eigenstate of $[\mathcal{C}]\mathcal{P}$. Outside of these special cases, the static-hadron-pair state in Eq. (5) is an equal-amplitude superposition of $\eta=+1$ and $\eta=-1$. The normalized projection onto each value of $\eta$ is
@@
$\Pi_{\eta}=\frac{\mathbbm{1}+\eta\,[\mathcal{C}]\mathcal{P}}{2}$ (10)
is the projector onto $[\mathcal{C}]\mathcal{P}$-parity $\eta$.
@@
with $J_{y}$ the generator of rotations around $\bm{\hat{y}}$. The action of $\mathcal{R}$ on the static-hadron-pair state in Eq. (5) can be obtained by applying the rotation operator $e^{-i\pi J_{y}}$ to both sides of Eq. (7e). The state on the right side of Eq. (7e) transforms as

### v3

# Model-independent predictions for decays of hidden-heavy hadrons into pairs of heavy hadrons

@@ Abstract
Hidden-heavy hadrons can decay into pairs of heavy hadrons through transitions from confining Born-Oppenheimer potentials to hadron-pair potentials with the same Born-Oppenheimer quantum numbers. The transitions are also constrained by conservation of angular momentum and parity. From these constraints, we derive model-independent selection rules for decays of hidden-heavy hadrons into pairs of heavy hadrons. The coupling potentials are expressed as sums of products of Born-Oppenheimer transition amplitudes and angular-momentum coefficients. If there is a single dominant Born-Oppenheimer transition amplitude, it factors out of the coupling potentials between hidden-heavy hadrons in the same Born-Oppenheimer multiplet and pairs of heavy hadrons in specific heavy-quark-spin-symmetry doublets. If furthermore the kinetic energies of the heavy hadrons are much larger than their spin splittings, we obtain analytic expressions for the relative partial decay rates in terms of Wigner 6-$j$ and 9-$j$ symbols. We consider in detail the decays of quarkonia and quarkonium hybrids into the lightest heavy-meson pairs. For quarkonia, our model-independent selection rules and relative partial decay rates agree with previous results from quark-pair-creation models in simple cases and give stronger results in other cases. For quarkonium hybrids, we find disagreement even in simple cases.
@@ I Introduction
The field of hidden-heavy-hadron spectroscopy began in 1974 with the discovery of the hidden-charm meson $J\!/\!\psi$ [1, 2]. By the end of the century, dozens of hidden-charm and hidden-bottom mesons had been discovered, all of which matched the pattern of quarkonium energy levels predicted by quark models. The spectrum of hidden-heavy hadrons was generally believed to be well understood. But all this changed in 2003 with the discovery of the $X(3872)$ meson [3]. Its observed decays revealed that it is a hidden-charm meson whose constituents include an additional light quark-antiquark pair. Since then, dozens of exotic hidden-charm and hidden-bottom hadrons have been discovered, inaugurating a renaissance of hidden-heavy-hadron spectroscopy; see, for instance, the reviews in Refs. [4, 5].
@@
In this paper, we focus on decays of hidden-heavy hadrons into pairs of heavy hadrons. These strong decays are subject to the low energy, nonperturbative regime of QCD, which is notoriously difficult to access from first principles. Hadron resonances with 2-body (and some 3-body) strong decay channels can be calculated ab initio using lattice QCD and the Lüscher formalism; see Ref. [6] for a review. But hidden-heavy hadrons that are above the threshold for a heavy-hadron pair have many multibody strong decay channels, including lower-energy hidden-heavy hadrons plus light hadrons. So, first-principles calculations of hidden-heavy hadron resonances using the Lüscher formalism might be impractical.
@@
An alternative approach to the study of hidden-heavy hadrons from first principles is the Born-Oppenheimer approximation for QCD [7]. The Born-Oppenheimer (B-O) approximation is based on the assumption that gluons and light quarks respond almost instantaneously to the motion of the heavy quark-antiquark pair, since the latter have a mass $m_{Q}$ that is much larger than the nonperturbative energy scale of QCD $\Lambda_{\textup{QCD}}$. In this approximation, the energy levels of the gluon and light-quark fields in the presence of static quark and antiquark sources, which can be calculated using lattice QCD, play the role of B-O potentials that determine the motion of the heavy quark-antiquark pair through a Schrödinger equation. It has been shown that the B-O approximation can be formulated as a rigorous effective theory of QCD [8, 9, 10, 11].
@@
There are two qualitatively different kinds of B-O potentials in QCD. The first kind consists of confining potentials that increase linearly at large distances [12]. The solutions to the Schrödinger equation in these potentials are bound states associated with compact hidden-heavy hadrons. The second kind consists of potentials that at large distances approach the scattering threshold of a pair of heavy hadrons [13]. The solutions to the Schrödinger equation in these potentials include scattering states of heavy-hadron pairs. They might also include bound states associated with heavy-hadron molecules.
@@
The B-O potentials for hidden-heavy hadrons and for heavy-hadron pairs mix with each other [13, 14]. Hence, any hidden-heavy hadron is associated with a pole of the S-matrix for heavy-hadron pairs. For hidden-heavy hadrons that are stable against decays into heavy-hadron pairs, the imaginary part of the pole is proportional to the total decay width into lower-energy hidden-heavy hadrons plus light hadrons. For hidden-heavy hadrons associated with heavy-hadron-pair scattering resonances, the imaginary part of the pole has additional contributions from decays into heavy-hadron pairs.
@@
One way to calculate partial decay widths of hidden-heavy hadrons into heavy-hadron pairs using the B-O approximation for QCD is to calculate self-energy corrections to “bare” hidden-heavy-hadron states by resummation of heavy-hadron loop diagrams, as was done in Refs. [15, 16]. But it should be noted that this approach can lead to inconsistent results for hidden-heavy hadrons that are strongly coupled to a nearby heavy-hadron-pair threshold [15]. Another way is to calculate resonances directly from the Schrödinger equation with coupled hidden-heavy-hadron and heavy-hadron-pair channels, as was done in Refs. [17, 18, 19]. This approach should be preferred as it respects unitarity and it yields consistent results for resonances that are close to a heavy-hadron-pair threshold [18].
@@
The remainder of this paper is organized as follows. In Section II, we define static hadrons and heavy hadrons. We then construct heavy-hadron-pair states and discuss their symmetries. In Section III, we derive general expressions for the coupling potentials between hidden-heavy hadrons and heavy-hadron pairs as sums of products of B-O transition amplitudes and angular-momentum coefficients. From these expressions, we derive model-independent selection rules for decays of hidden-heavy hadrons. In some cases, we also obtain analytic predictions for relative partial decay rates. In Section IV, we discuss in detail the decays of quarkonia and quarkonium hybrids into pairs of heavy mesons. In Section V, we compare our results with previous ones from quark-pair-creation models. Finally, we summarize these results in Section VI.
@@ II.1 Single Heavy Hadron
$\mathcal{P}\biggl{|}\begin{matrix}\frac{1}{2}^{\pm}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ =±$=\pm\biggl{|}\begin{matrix}\frac{1}{2}^{\pm}\\[-5.0pt] m\end{matrix}\biggr{\rangle},$ (1a)
$\mathcal{C}\biggl{|}\begin{matrix}\frac{1}{2}^{\pm}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ $=\biggl{|}\begin{matrix}\frac{1}{2}^{\mp}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (1b)
@@
$\mathcal{P}\biggl{|}\begin{matrix}j^{\pi}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ $=\pi\biggl{|}\begin{matrix}j^{\pi}\\[-5.0pt] m\end{matrix}\biggr{\rangle},$ (2a)
$\mathcal{C}\biggl{|}\begin{matrix}j^{\pi}\\[-5.0pt] m\end{matrix}\biggr{\rangle}$ $=\biggl{|}\begin{matrix}j^{\pi(-1)^{2j}}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (2b)
Note that a static hadron and its charge conjugate have the same (opposite) parity if $j$ is an integer (half-integer), hence the multiplicative factor $(-1)^{2j}$ in the parity superscript on the right side of Eq. (2b).
@@
where $j$ can be $J+\frac{1}{2}$ or $\bigl{\lvert}J-\frac{1}{2}\bigr{\rvert}$ and $\bigl{\langle}\begin{smallmatrix}j_{1}&j_{2}\\ m_{1}&m_{2}\\ \end{smallmatrix}\big{|}\begin{smallmatrix}j_{3}\\ m_{3}\\ \end{smallmatrix}\bigr{\rangle}$ is a Clebsch-Gordan coefficient. The hadron is a meson if $J$ is an integer and a baryon if $J$ is a half-integer. We have adopted the standard order of writing the spin/parity of the heavy quark or antiquark first and the angular-momentum/parity of the light-QCD fields second. From Eqs. (1)-(2), the transformations under $\mathcal{P}$ and $\mathcal{C}$ of the heavy-hadron state in Eq. (3) are:
$\mathcal{P}\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j^{\pi}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle}$ =±$=\pm\pi\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j^{\pi}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle},$ (4a)
$\mathcal{C}\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j^{\pi}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle}$ $=\biggl{|}\Bigl{(}\tfrac{1}{2}^{\mp},j^{\pi(-1)^{2j}}\Bigr{)}\begin{matrix}J\\[-5.0pt] M\end{matrix}\biggr{\rangle}.$ (4b)
@@
Since the spin of a heavy quark or antiquark is conserved up to corrections suppressed by $1/m_{Q}$, the mass difference between two heavy hadrons corresponding to the same static hadron $j^{\pi}$ but with different spins $J=\bigl{\lvert}j-\frac{1}{2}\bigr{\rvert}$ and $J=j+\frac{1}{2}$ must also be suppressed by $1/m_{Q}$. Therefore, heavy hadrons form approximately degenerate doublets of heavy quark spin symmetry (HQSS) labeled by the light-QCD angular-momentum/parity $j^{\pi}$ and other quantum numbers.
@@ II.2 Pair of Static Hadrons
Let us now consider a pair of static hadrons at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$ with light-QCD quantum numbers $j_{1}^{\pi_{1}}$ and $j_{2}^{\pi_{2}}$, respectively. We take the first static hadron to contain the $\bm{3}$ source and the second to contain the $\bm{3}^{\ast}$ source. Since the angular-momentum/parity quantum numbers correspond to rotations and reflections around different points in space, we specify the position ±$\pm\frac{1}{2}\bm{r}$ by an argument (±)plus-or-minus($\pm$)( ± ). The static-hadron-pair states are therefore labeled by $\bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\bigr{)}$.
@@
Light QCD in the presence of static $\bm{3}$ and $\bm{3}^{\ast}$ sources at $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$ has cylindrical symmetries consisting of rotations around the $\bm{\hat{r}}$ axis and reflections through planes containing $\bm{\hat{r}}$. We denote the reflection corresponding to a specific plane (which need not be specified) by $\mathcal{R}$. It also has a discrete symmetry under $\mathcal{C}\mathcal{P}$, the combination of charge-conjugation and parity. We refer to the group formed by the cylindrical symmetries and the $\mathcal{C}\mathcal{P}$ symmetry as the B-O symmetry group.
@@
The B-O symmetries imply that the eigenstates of the light-QCD Hamiltonian can be chosen to be simultaneously eigenstates of $\bm{J}_{\textup{light}}\cdot\bm{\hat{r}}$, that is, the projection of the light-QCD angular momentum onto the axis passing through the sources. We denote its eigenvalues by $\lambda$. They can be also chosen to be simultaneously eigenstates of $\mathcal{C}\mathcal{P}$. We denote its eigenvalues by $\eta$. In the sector with $\lambda=0$, the reflection operator $\mathcal{R}$ can also be diagonalized. We denote its eigenvalues by $\epsilon$. We refer to $\lambda$ and $\eta$ and also $\epsilon$ if $\lambda=0$ as B-O quantum numbers. Alternatively, the light-QCD Hamiltonian, $\lvert\bm{J}_{\textup{light}}\cdot\hat{\bm{r}}\rvert$, $\mathcal{C}\mathcal{P}$, and $\mathcal{R}$ can be simultaneously diagonalized. Thus an alternative choice for the B-O quantum numbers is $\lvert\lambda\rvert$, $\eta$, and $\epsilon$. It is customary to denote these B-O quantum numbers using the notation $\Lambda_{\eta}^{\epsilon}$, where $\Lambda=\lvert\lambda\rvert$, the subscript $\eta$ is $g$ or $u$ if the $\mathcal{C}\mathcal{P}$ eigenvalue $\eta$ is $+1$ or $-1$, and the superscript $\epsilon$ is $+$ or -- if the reflection eigenvalue $\epsilon$ is $+1$ or $-1$. If $\Lambda>0$, the superscript $\epsilon$ is often omitted because cylindrical symmetry requires the $\Lambda_{\eta}^{+}$ and $\Lambda_{\eta}^{-}$ states to be degenerate in energy. It is commonplace to specify integer values of $\Lambda$ with an uppercase Greek letter instead of a number, according to the code $\Lambda\to\Sigma,\Pi,\Delta$ for $\Lambda=0,1,2$ and so on.
@@
The sign is the product of a factor $(-1)^{j_{1}+j_{2}-j^{\prime}}$ from the symmetries of Clebsch-Gordan coefficients and a factor $(-1)^{4j_{1}j_{2}}$ from changing the order of the light-QCD operators. [^1] The factor $(-1)^{4j_{1}j_{2}}$ is $-1$ if both static hadrons are static mesons and $+1$ otherwise. From Eqs. (2), the transformations under $\mathcal{P}$ and $\mathcal{C}$ of the static-hadron-pair state in Eq. (5) are:
$\mathcal{P}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}=$ $\pi_{1}\pi_{2}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(-),j_{2}^{\pi_{2}}(+)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle},$ (7a)
$\mathcal{C}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}=$ $\biggl{|}\Bigl{(}j_{1}^{\pi_{1}(-1)^{2j_{1}}}(+),j_{2}^{\pi_{2}(-1)^{2j_{2}}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}.$ (7b)
On the right side of Eq. (7a), the position labels ()(-)( - ) and (+)(+)( + ) can be put into the standard order by using Eq. (6).
The static-hadron-pair state in Eq. (5) is a simultaneous eigenstate of $\mathcal{C}\mathcal{P}$ when the two static hadrons are charge conjugates, which requires $j_{2}^{\pi_{2}}=j_{1}^{\pi_{1}(-1)^{2j_{1}}}$. After reordering the ()(-)( - ) and (+)(+)( + ) states using Eq. (6), the $\mathcal{C}\mathcal{P}$ eigenvalue is the product of $\pi_{1}^{2}(-1)^{2j_{1}}=(-1)^{2j_{1}}$ and $(-1)^{4j_{1}^{2}+2j_{1}-j^{\prime}}$. Since $2j_{1}(2j_{1}+1)$ is even, the $\mathcal{C}\mathcal{P}$ eigenvalue reduces to
$\eta=(-1)^{j^{\prime}+2j_{1}},$ (8)
where we have used the fact that $j^{\prime}$ is an integer since $j_{1}=j_{2}$. Outside of this special case, the static-hadron-pair state in Eq. (5) is an equal-amplitude superposition of $\eta=+1$ and $\eta=-1$. The normalized projection onto each value of $\eta$ is
@@
$\Pi_{\eta}=\frac{\mathbbm{1}+\eta\,\mathcal{C}\mathcal{P}}{2}$ (10)
is the projector onto $\mathcal{C}\mathcal{P}$-parity $\eta$.
@@
with $J_{y}$ the generator of rotations around $\bm{\hat{y}}$. The action of $\mathcal{R}$ on the static-hadron-pair state in Eq. (5) can be obtained by applying the rotation operator $e^{-i\pi J_{y}}$ to both sides of Eq. (7a). The state on the right side of Eq. (7a) transforms as